OpenStudy (anonymous):

first 7 + x - 15 = -2 2/3 second 18 - 3n + 2 = n + 20 - 4n? help please

OpenStudy (anonymous):

7+x-15 = -8/3 x-8 = -8/3 x = 8-8/3 x = 16/3

OpenStudy (anonymous):

7-15+x=-22/3 -8+x=-22/3 x=-22/3+8 x=-22/3+24/3 x=2/3 second 18-3n+2=n-4n+20 20-3n=-3n+20 -3n+3n=20-20 0=0 i think its wrong...so pls check

OpenStudy (anonymous):

18 - 3n + 2 = n + 20 - 4n 20-3n = 20 - 3n 0 = 0 :D

OpenStudy (anonymous):

babs is it 22/3 or 2(2/3)?

OpenStudy (anonymous):

i think he represented -2 2/3 as a mixed fraction :/

OpenStudy (anonymous):

humm i donno..if its a mixed fraction urs is gud otherwise mine is gud...:)

OpenStudy (anonymous):

the first one u guys gave me isnt on there

OpenStudy (anonymous):

here is what is on there for the answers 5 1/3 6 1/3 10 2/3 20 2/3